# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v10

@@ 6.2 Neighborhood Density
The power of the Decreasing Neighborhood Sequence Property (DNSP) and the GLOVER data structure goes well beyond the above two lemmas that have been mentioned in literature in some form, as we will prove. In fact, the above lemmas can be generalized, but first we need to show that the DNSP can show more about the overall graph. Under the standing assumption that every node satisfies DNSP, $|N^{++}(u)|<|N^{+}(u)|\forall u\in V(G)$, we proceed to explore Neighborhood Density. This concept will extend beyond pairwise neighborhood comparisons to offer insights into the broader structural constraints imposed by DNSP on the graph.
@@
The earlier lemmas (Interior and Exterior Load Balancing) are specific instances of how DNSP restricts node neighborhoods locally. However, DNSP’s implications reach further, shaping the overall topology and connectivity patterns—especially when analyzed in the context of the rooted neighborhoods and the GLOVER data structure.
@@
One additional thing that makes the rooted neighborhoods important is that they are each their own instance of the SSNC. What we mean by that is that we can use a divide and conquer approach, and individually isolate the rooted neighborhood $R_{i}$ from the rest of the graph. The SSNC would apply to the rooted neighborhood $R_{i}$. This rooted neighborhood $R_{i}$ is also an oriented graph and so it can be put into a GLOVER data structure. This gives a relationship between the nodes in the rooted neighborhoods and different ways they can interact with nodes outside the rooted neighborhoods. Because these nodes have been built within this data structure, we do not need to understand them in terms of any oriented graph. Instead, we have more of a question of the preventative measures we have been taking to avoid degree doubling nodes.
@@
We have discussed how nodes are required to do load balancing to support parent nodes. What nodes also need to do is have their interior degrees double. What we mean by this is that, in the absence of back arcs, a node’s degree equation can be split into interior and exterior degrees. In the rooted neighborhood $R_{i}$, the node $u_{i}$ will then have the out-degree of $\text{int}(u_{i-1},u_{i})$, where $u_{i-1}\in R_{i-1}$ is a parent of $u_{i}$. The statement that the interior degree doubles is just saying that the SSNC holds for the node $u_{i}$ when looking at the interior degree as its out-degree in $R_{i}$.
@@
This gives us a way backward and a way forward. The path backward is the nodes in a current neighborhood $R_{i}$ use load balancing to help their parents to avoid becoming degree doubling nodes. The way forward is through interior degree doubling. With this, a node’s degree equation, which was reduced to interior degree and exterior degree, is simply focused on exterior degree going forward. Once we have proven that the interior degree doubles, we will be able to simply search for nodes whose exterior degree doubles as well. This will open the door for our path-finding approach through exterior nodes.
@@
Before we discuss the interior degree doubling lemma, let’s introduce it with an example. We have seen with Lemma 6.2 that we need to have a cycle in the first rooted neighborhood, but this does not tell us anything about the remaining neighborhoods. The next Lemma, 6.3 lets us know that for every other node the exterior degree must be less than the degree of that node. This gives us a relationship going forward about the decreasing nature of the rooted neighborhoods.
@@
Our assumption in this example is that all nodes have the DNSP. We begin with the node $v_{0}$. Because of the DNSP, we would like to have every node $u\in N^{+}(v+0)$ to have interior degrees of at least 1. Similar to Example 4.2, this cycle in the rooted neighborhood $R_{1}$ allows the nodes in $R_{1}$ to have less than 7 exterior neighbors, while still having their required number of out-neighbors. These exterior neighbors for the nodes in $R_{1}$ are the second neighbors for $v_{0}$. Since there are less than 6 nodes in $R_{2}$, we see that $v_{0}$ is not a degree doubling node, since 6 ¡ 7.
Continuing on to the nodes in $R_{1}$, they also have minimum out-degree 7 when we consider their interior and exterior degrees. Every node in $R_{2}$ has an interior degree of 2. Thus in order for nodes in $R_{2}$ to meet the minimum out-degree requirements, there must be at least 5 nodes in $R_{3}$. Again, these two cycles in $R_{2}$ help prevent the degree doubling of any node in $R_{1}$. Each node in $R_{3}$ is a second neighbor of a node in $R_{1}$. Because there are 5 nodes in $R_{3}$, we can show that that no node in $R_{1}$ is a degree-doubling node. We can say this, first, because the interior degrees double due to cycles, but next because the exterior degrees do not double because 5 ¡ 6.
@@
Next, we move to the nodes in the rooted neighborhood $R_{2}$. What we first consider is the case that some node $v_{2}\in R_{2}$ has an interior degree less than 3. This implies that $v_{2}$ will have at least two five exterior neighbors. This means that a parent $u_{1}\in R_{1}$ of $v_{2}$ will have five exterior neighbors first neighbors and five exterior second neighbors, causing $u_{1}$’s degree to double. This means that we need every node in $R_{2}$ to have interior degree 3.
@@
Unfortunately, what we see is that the rooted neighborhood $R_{3}$ requires an oriented graph of 5 nodes, each of degree 3 inside it to support load balancing. This is not possible to construct. To see this, the complete simple graph on 5 nodes has 10 edges. However, every node with out-degree 3 requires 15 arcs. So we are requiring more arcs than is possible in an orientation of the complete graph of 5 vertices. This means that some node $u\in R_{3}$ will have an exterior degree at least 5. Because of this, we have a violation of the DNSP. By the cycle constructions as we have previously seen, we will be able to have interior degrees doubling, so this violation of the DNSP leads to a node in $R_{2}$ becoming a degree doubling node.
@@
Example 6.2 just showed a few things. The first is how the minimum degree 7 case might look in a GLOVER data structure. A few assumptions had to be made for simplicity; for example, we illustrated a rooted neighborhood $R_{1}$ assuming its nodes have an interior degree of exactly one rather than at least one. This assumption allowed us to see that every node within that neighborhood becomes a part of an interior cycle. This reinforces our strategy of splitting a node’s total degree into its interior and exterior components. Furthermore, we observe that because the interior degree is part of a cycle, every node within that cycle will have its interior degree double (i.e., its second interior neighborhood will be at least as large as its first). As a result of this, if we can prove that the exterior degree will also consistently be such that the total second neighborhood of a node is greater than or equal to its total first neighborhood, then we will have an algorithm for the Seymour Second Neighborhood Conjecture (SSNC).
@@
Our next lemma generalizes this example. We will show that our strategy of first dividing a node’s total degree into its interior degree and exterior degree, and then demonstrating that the interior degree will always double (i.e., its second interior neighborhood will be at least as large as its first), allows us to gain significant ground on the problem. This approach effectively separates the problem into analyzing individual rooted neighborhoods, which are essentially local instances of the Seymour Second Neighborhood Conjecture. Each rooted neighborhood $R_{i}$ is independent of other rooted neighborhoods $R_{j}$, requiring only the information contained within its own nodes and arcs for internal analysis. This is similar to a divide and conquer approach, but without recursion. Knowing that a node’s interior degree has doubled means we only need to examine its exterior neighbors to determine if that node’s overall second neighborhood is at least as large as its first, which is the core of the conjecture.
@@
Interior Degree Doubling
Let $G$ be an oriented graph, and let $v_{0}$ be a node of minimum out-degree in $G$. Then, for every rooted neighborhood $R_{i}$ at distance $i$, there exists at least one node whose second interior out-degree is at least as large as its first interior out-degree:
@@
We provide a constructive algorithm and use properties of permutation groups to show that the condition holds. The proof relies on the fact that the internal structure of $R_{i}$ is already known to be a permutation into disjoint cycles of length $\geq 3$
This proof relies on the established property (as proven in Section 3) that for the class of oriented graphs under consideration, every rooted neighborhood $R_{i}$ forms an induced subgraph consisting of a permutation of its nodes, decomposing into disjoint cycles, each of length at least 3. This inherent structure ensures that within $R_{i}$, each node has exactly one outgoing and one incoming edge.
For a node $v\in R_{i}$, its first interior out-degree, denoted $|N^{+}_{int}(v)|$, refers to the number of edges $(v,u)$ such that $u\in R_{i}$. Similarly, the second interior out-degree of $v$, denoted $|N^{++}_{int}(v)|$, refers to the number of nodes $w\in R_{i}$ such that there exists a node $u\in R_{i}$ with edges $(v,u)$ and $(u,w)$
We begin this proof with a statement of a simple algorithm and will show that this algorithm is feasible for the Interior Degree Doubling. The purpose of this algorithm is to give an example as a generalization of a load balancing algorithm, and show this is possible. This is not the only such algorithm. Indeed, the number of derangements $D_{n}$ grows at a complexity of $\frac{n!}{e}$. That fraction includes transpositions, or 2-cycles which are outlawed. Removing the number of 2-cycles from this growth rate does not reduce this by much.
We outlawed 2-cycles because they are not allowed in oriented graphs and we would like this permutation to transpose back to our graph theoretical framework. To that nature, we have already proved in Section 3 (Initial Lemmas) that all our counter-examples must have minimum out-degree at least 3. This agrees with the minimum degree requirements for a cycle.
Algorithm (MapInteriorDegrees):
  • Input: A list of nodes $R_{i}$, and integer distance $i$
  • Goal: Assign each node in $R_{i}$ exactly $i$ interior neighbors
  • Initialize a dictionary $assigned\_neighbors$ such that each node maps to an empty set
  • Let $n\leftarrow|R_{i}|$
  • For each node $u\in R_{i}$:
    • Set $assigned\leftarrow 0$
    • For $step=1$ to $n-1$:
        Let $v\leftarrow R_{i}[(index(u)+step)\mod n]$
        If $v\neq u$ and $v\notin assigned\_neighbors[u]$:
          Add $v$ to $assigned\_neighbors[u]$
          Increment $assigned\leftarrow assigned+1$
        If $assigned=i$, break
  • Output: The mapping $assigned\_neighbors$, where each node has exactly $i$ interior neighbors
Figure 11: Given a rooted neighborhood $R_{i}$, this algorithm assigns each node to participate in exactly $i$ interior cycles, simulating repeated interior degree mappings for each level. This construction ensures every node maps to interior neighbors via multiple cycle layers, supporting the interior degree doubling condition.
@@
Each node appears once in each cycle. These two permutations correspond to two independent interior mappings. Since every node has a first interior neighbor in each cycle, and each of those has its own successor, this yields:
Even with limitations on 2-cycles, the intuition behind this algorithm will allow for the utilization of combinatorial space in $R_{i}$. We can actually look at this as we begin in the rooted neighborhood $R_{i}$.
@@
This results in $i!$ possible candidate paths (modulo orientation constraints). This number of such derangements still grows factorially at roughly $\frac{n!}{\exp}$. This indicates that we still have an abundance of cycle permutations with the 2-cycle restrictions.
The structure of these rooted neighborhoods forms a permutation on the set $R_{i}$, which decomposes into disjoint cycles. Since oriented graphs prohibit 2-cycles, each of these cycles must have a length $k\geq 3$.A key result from group theory (Lagrange’s Theorem) tells us that the order of any subgroup divides the order of the full symmetric group on $R_{i}$, $S_{|R_{i}|}$. These permutations decompose into disjoint cycles whose lengths divide $|R_{i}|$, and the cycles are structure-preserving under composition.
Each node is mapped to a unique successor in its cycle. Therefore, in every cycle of length $k\geq 3$, each node has out-degree 1 and maps to another node in the same cycle, producing a single out-edge within $R_{i}$. The node it maps to will also have a unique out-neighbor (since cycles do not repeat nodes), yielding a second-level interior neighbor.
Consider any node $v$ within one of these cycles of length $k\geq 3$. By definition of a cycle within $R_{i}$, $v$ has exactly one outgoing edge $(v,u)$ where $u\in R_{i}$. Therefore, $|N^{+}_{int}(v)|=1$. Furthermore, since $k\geq 3$, $u$ must also have exactly one outgoing edge $(u,w)$ where $w\in R_{i}$ and $w\neq v$. Thus, $w$ is a unique second-level interior neighbor, implying $|N^{++}_{int}(v)|=1$.
Since for every node $v$ in any such rooted neighborhood $R_{i}$, we have $|N^{+}_{int}(v)|=1$ and $|N^{++}_{int}(v)|=1$, the condition $|N^{++}_{int}(v)|\geq|N^{+}_{int}(v)|$ is satisfied $(1\geq 1)$. As this holds for every node in $R_{i}$, it trivially holds for ’at least one node’.
@@
This result is also aligned with Lemma 4.1 (Minimum Out-Degree $<3$) and related results from [10], confirming that low out-degree nodes within cycles tend to satisfy the degree doubling property.
We should note that both single and multiple cycle decompositions will satisfy the degree doubling requirement, as all meet the requirement of having all nodes with their interior second degree being at least that of their interior first degree. So whether we use the MapInteriorDegrees algorithm to construct either of these, the result will still be interior degrees doubling.
@@
The positive answer to interior degree-doubling is helpful to the construction of our algorithm. In the absence of back arcs, we can split a node’s degree into interior and exterior degrees. Knowing then that a node’s interior degree is doubled in the rooted neighborhood $R_{i}$ means that if that node’s exterior degree is also doubled, then that node’s overall degree will double. This gives light to our overall strategy. We will assume that all nodes are bound by the Decreasing Neighborhood Sequence Property (DNSP). Then if any node has both its interior and exterior degree double, its overall degree will double, which will lead to a contradiction of the DNSP.
@@
Lemma 6.4 (Interior Degrees Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the rooted neighborhoods as an oriented graph and searched for a node whose degree doubled inside that rooted neighborhood. The proof uses abstract algebra to show that the degree of a node that is necessary to take place inside a rooted neighborhood will not be arbitrary. Instead, these will be cycles, as we have been showing with our load balancing examples. With these cycles, it is easy to show that every node on every cycle will have its interior degree double because it has out-degrees of 1.
@@
This also shows the reach that SSNC has across mathematics. Among the published literature there is not much work in the field of abstract algebra connected to the SSNC. However, given the importance of cycles in reducing keeping nodes from having their degrees double, we are able to see that we are always able to find a representation that maps nodes to in
@@
By finding a positive answer to this localized version of the conjecture, and not just a positive answer but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture. Not only that, but it helps us to dissect the problem down into two components: interior degree and exterior degree. We know that the interior degree of every node inside the rooted neighborhoods doubles, so we can turn our attention to the exterior degrees.
@@
We are then able to introduce a traversal algorithm that explores these exterior degrees. It will systematically search the nodes of $G$ for a node whose exterior degree doubles. That is, whose second (exterior) neighborhood is at least as large as its first. Lemma 6.5 (Decreasing Exteriors Lemma) will show that these exterior neighborhoods must shrink at each step. Once we have found a node that satisfies this doubling condition, we will have found our contradiction.
We give particular note the difference between this data structure and other well known data structures. While this data structure will have similarity to other computer science and graph theory structures, those structures were not constructed for this conjecture. As such, they do not decipher the different types of arcs in a manner towards proving this conjecture. If traversal algorithms like Breadth First Search or Depth First Search could solve this problem, this would have been solved long ago. Instead, this data structure needs to be able to build a traversal algorithm that can do the things that BFS and DFS could not do with respect to this conjecture. In particular, the traversal portion of the algorithm we are constructing does not begin until we know that the interior degrees within each rooted neighborhood double. This helps apply the deconstruction of a node’s degree into interior, exterior and back arc degrees. Once we can show that back arcs do not hurt the algorithm, we can continue with a path traversal algorithm through exteriors.

### v12

@@ 6.2 Neighborhood Density
The power of the Decreasing Neighborhood Sequence Property (DNSP) and the Graph Level Order goes well beyond the above two lemmas that have been mentioned in literature in some form, as we will prove. In fact, the above lemmas can be generalized, but first we need to show that the DNSP can show more about the overall graph. Under the standing assumption that every node satisfies DNSP, $|N^{++}(u)|<|N^{+}(u)|\forall u\in G$, we proceed to explore Neighborhood Density. This concept will extend beyond pairwise neighborhood comparisons to offer insights into the broader structural constraints imposed by DNSP on the graph.
@@
The earlier lemmas (Interior and Exterior Load Balancing) are specific instances of how DNSP restricts node neighborhoods locally. However, DNSP’s implications reach further, shaping the overall topology and connectivity patterns—especially when analyzed in the context of the rooted neighborhoods and the Graph Level Order.
@@
One additional thing that makes the rooted neighborhoods important is that they are each their own instance of the SSNC. What we mean by that is that we can use a divide and conquer approach, and individually isolate the rooted neighborhood $R_{i}$ from the rest of the graph. The SSNC would apply to the rooted neighborhood $R_{i}$. This rooted neighborhood $R_{i}$ is also an oriented graph and so it can be put into a Graph Level Order. This gives a relationship between the nodes in the rooted neighborhoods and different ways they can interact with nodes outside the rooted neighborhoods. Because these nodes have been built within this data structure, we do not need to understand them in terms of any oriented graph. Instead, we have more of a question of the preventative measures we have been taking to avoid degree doubling nodes.
@@
We have discussed how nodes are required to do load balancing to support parent nodes. What nodes also need to do is have their ’interior’ degrees double. What we mean by this is that, in the absence of back arcs, we know that a node’s degree equation can be split into interior and exterior degrees. In the rooted neighborhood $R_{i}$, the node $u_{i}$ will then have the out-degree of $\text{int}(u_{i-1},u_{i})$, where $u_{i-1}\in R_{i-1}$ is a parent of $u_{i}$. The statement that the interior degree doubles is just saying that the SSNC holds for the node $u_{i}$ when looking at the interior degree as its out-degree in $R_{i}$.
@@
This gives us a way backward and a way forward. The path backward is the nodes in a current neighborhood $R_{i}$ using load balancing to help their parents to avoid becoming degree doubling nodes. The way forward is through interior degree doubling. With this, a node’s degree equation, which was reduced to interior degree and exterior degree, is simply focused on exterior degree going forward. Once we have proven that the interior degree doubles, we will be able to simply search for nodes whose exterior degree doubles as well. This will open the door for our path-finding traversal through exterior nodes.
@@
Before we discuss the interior degree doubling lemma, let’s introduce it with an example. We have seen with Lemma 6.2 (Interior Load Balancing) that we need to have a cycle in the first rooted neighborhood, but this does not tell us anything about the remaining neighborhoods. The next Lemma, 6.3 (Exterior Load Balancing)lets us know that for every other node the exterior degree must be less than the degree of that node. This gives us a relationship going forward about the decreasing nature of the rooted neighborhoods.
@@
Our assumption in this example is that all nodes have the DNSP. We begin with the node $v_{0}$. Because of the DNSP and Lemma 6.2 (Interior Load Balancing), know that every node $u\in R_{1}$ will have $\text{int}(v_{0},u)\geq 1$. Similar to Example 4.2, the cycle(s) in the rooted neighborhood $R_{1}$ allows the nodes in $R_{1}$ to have less than 7 exterior neighbors, while still having their required number of out-neighbors. These exterior neighbors for the nodes in $R_{1}$ are the second neighbors for $v_{0}$. Since there are less than 6 nodes in $R_{2}$, we see that $v_{0}$ is not a degree doubling node, since $6<7$. The situation where every node in $R_{1}$ connects to more than one interior arc means that the exterior degree of every node in $R_{1}$ simply needs to be less than 6 (Every node in $R_{1}$ having 2 interior arcs and 5 exterior arcs would suffice, every node in $R_{1}$ having 3 interior arcs having 3 interior arcs and 4 exterior arcs would suffice as well. Both these situations lead to the sequence simply shrinking faster. As there are only $\binom{7}{2}=21$ arcs, nodes cannot have an average degree higher than 3. ). No matter the situation with the remaining arcs, the cycle in $R_{1}$ is what allows $v_{0}$ to not be a degree doubling node.
@@
Continuing on to the nodes in $R_{2}$, they also have a minimum overall out-degree of at least 7. We want to show that every node in $R_{2}$ will need to have an interior out-degree of at least 2. To see this, consider their exterior neighbors, which map to the rooted neighborhood $R_{3}$. By the definition of exterior neighbors, the exterior neighbors in $R_{3}$ are second neighbors of parent nodes in $R_{1}$ and first neighbors of sibling nodes in $R_{2}$. If a node in $u_{2}\in R_{2}$ has only one interior neighbor in $R_{2}$, then $u_{2}$ must have at least six exterior neighbors in $R_{3}$.
Remember Lemma 5.1 (Partition of Node’s Degree) that states that a node’s degree can be partitioned into its interior out-degree, exterior out-degree and back arcs. We are under the assumption of no back arcs. Next, we notice that for nodes of low degree, as a parent $u_{1}\in R_{1}$ of $u_{2}$, they have their degrees double trivially [20]. Combine this with the fact that this same parent’s exterior degree will now double by exterior degree traversal, because two neighborhoods are the same size $|R_{2}|=|R_{3}|$. That is $|R_{2}|$, the number of first exterior neighbors of $u_{1}$, and $|R_{3}|$, the number of second exterior neighbors of $u_{1}$ are equal in size. By this equality we have that $u_{1}$’s exterior degree doubles. We already showed that $u_{1}$’s interior degree doubles, and since we are assuming that there are no back arcs we have that $u_{1}$’s overall degree doubles. So a node in $R_{2}$ cannot have an interior out-degree of $<2$. This implies that every node in $R_{2}$ must have interior out-degree of at least 2.
Consider the case where a node $u_{2}\in R_{2}$ has a higher interior out-degree than 2. The most balanced degree distribution of the $\binom{6}{2}=15$ possible arcs is (3, 3, 3, 2, 2, 2). Even in this distribution we see that there exists a node with interior out-degree 2, so $R_{3}$ would need to have 5 nodes. More skewed distributions lead to even more nodes with interior out-degree of 2 or less. Consider the distributions (4, 3, 2, 2, 2, 2) and (5, 2, 2, 2, 2, 2). They both lead to even more nodes of degree two. Thus in order to meet the minimum out-degree requirements, there must be at least 5 nodes in $R_{3}$. Five is a prime number, so we know that every node in $R_{2}$ will map to two disjoint cycles in $R_{3}$ of length five. Like in Lemma 6.2 (Interior Load Balancing), these two cycles in $R_{3}$ help prevent the degree doubling of any node in $R_{2}$. Each node in $R_{3}$ is a second neighbor of a node in $R_{1}$. Let us perform calculations on the nodes in $R_{1}$. First, we have already calculated their interior degrees, and because of their low degree, we know that every node in $R_{1}$ will have their degree in $R_{1}$ double. Next, we calculate, their first and second exterior degrees. The first exterior degree of the node $u_{1}\in R_{1}$ will be $|R_{2}|=6$. The second exterior degree of the node $u_{1}\in R_{1}$ will be $|R_{3}|=5$. Because we have $|R_{2}|>|R_{3}|$, the node $u_{1}$ has more first neighbors than second neighbors and $u_{1}$ is not a degree doubling node. The same is true for every node in $R_{1}$.
Next, we move to the nodes in the rooted neighborhood $R_{3}$. Consider, first, the case that some node $v_{3}\in R_{3}$ has an interior degree less than 3. Similar to the $R_{2}$ case, we know that the node $v_{3}$ has a parent $v_{2}\in R_{2}$ that has interior degree of two. What we were able to show by Lemma 4.1 (Minimum Out-Degree $<$ 3) is that nodes with these small degrees will be degree doubling nodes. This means that $v_{2}$ will have its interior degree double. We need to consider the exterior degree of $v_{2}$. Since $\text{int}(v_{2},v_{3})\leq 2$, we have that $\text{ext}(v_{2},v_{3})\geq 5$. By definition, the exterior neighbors that $v_{2}$ has through $v_{3}$ are the second neighbors of $v_{2}$ that are first neighbors of $v_{3}$. This places $\text{ext}(v_{2},v_{3})$ into the next rooted neighborhood $R_{4}$ since every node in $R_{2}$ is a parent of every node in $R_{3}$ by our construction. This implies that $v_{3}$ will have at least five exterior neighbors. This means that $v_{2}\in R_{2}$ of will have five first exterior neighbors and five exterior second neighbors, causing it’s degree to double, which we do not want. This means that we need every node in $R_{3}$ to have interior out-degree 3.
@@
Unfortunately, what we see is that the rooted neighborhood $R_{3}$ requires an oriented graph of 5 nodes, each of degree 3 inside it to support load balancing. This is not possible to construct. To see this, the complete simple graph on 5 nodes has 10 arcs. However, every node with out-degree 3 requires $5\cdot 3=15$ arcs. Thus, we are requiring more arcs than is possible in an orientation of the complete graph of 5 vertices. This means that some node $u\in R_{3}$ will have an exterior degree at least 5. Because of this, we have a violation of the DNSP. By the cycle constructions as we have previously seen, we will be able to have interior degrees doubling, so this violation of the DNSP leads to a node in $R_{2}$ becoming a degree doubling node.
@@
Example 10 just showed how the minimum degree 7 case looks in a Graph Level Order. We also saw how beneficial it is to our strategy of splitting a node’s total degree into its interior and exterior components. Furthermore, we observe that because the interior degree is part of a cycle within the first neighborhood, every node within that cycle will have its interior degree doubled. As a result of this, if we can prove that the exterior degree will also consistently be such that the total second neighborhood of a node is greater than or equal to its total first neighborhood, then we will have an algorithm for the SSNC.
@@
Our next lemma generalizes this example. We will show that our strategy of first dividing a node’s total degree into its interior degree and exterior degree, and then demonstrating that the interior degree will always double, allows us to gain significant ground on the problem. This approach effectively separates the problem into analyzing individual rooted neighborhoods, which are essentially local instances of the (SSNC). Each rooted neighborhood $R_{i}$ is independent of other rooted neighborhoods $R_{j}$, requiring only the information contained within its own nodes and arcs for internal analysis. This is similar to a divide and conquer approach, but without recursion. Knowing that a node’s interior degree has doubled means we only need to examine its exterior neighbors to determine if that node’s overall second neighborhood is at least as large as its first, which is the core of the conjecture.
@@
Interior Degree Doubling Let $G$ be an oriented graph, and let $v_{0}$ be a node of minimum out-degree in $G$. If nodes at distance $i$ must have degree at least $i)$, then for every rooted neighborhood $R_{i}$ at distance $i$, there exists at least one node whose second interior out-degree is at least as large as its first interior out-degree:
@@
We will provide a constructive algorithm and use properties of permutation groups to show that the condition holds. That is, we will show that we can always map each node in $R_{i}$ to $i$ disjoint cycles. The proof relies on the fact that the internal structure of $R_{i}$ is already known to be a permutation into disjoint cycles of length $\geq 3$. This inherent structure ensures that within $R_{i}$, each node has exactly one outgoing and one incoming edge.
For a node $v\in R_{i}$, its first interior out-degree, denoted $|N^{+}_{\text{int}}(v)|$, refers to the number of nodes $v\to u$ such that $u\in R_{i}$. Similarly, the second interior out-degree of $v$, denoted $|N^{++}_{\text{int}}(v)|$, refers to the number of nodes $w\in R_{i}$ such that there exists a node $u\in R_{i}$ with edges $v\to u\to w$.
We begin this proof with a statement of a simple algorithm and will show that this algorithm is feasible for the Interior Degree Doubling. The purpose of this algorithm is to give an example as a generalization of a load balancing algorithm, and show this is possible. This is not the only such algorithm. Indeed, the number of derangements $D_{n}$ grows at a complexity of $\frac{n!}{e}$. That fraction includes transpositions, or 2-cycles which are outlawed. Removing the number of 2-cycles from this growth rate does not reduce this by much. We outlaw 2-cycles because they are not allowed in oriented graphs and we would like this permutation to transpose back to our graph theoretical framework. To that nature, we have already proved in Section 4 (Initial Lemmas) that all our counter-examples must have minimum out-degree at least 3. This agrees with the minimum degree requirements for a cycle.
Algorithm 1 MapInteriorDegrees
1:A list of nodes $R_{i}$, integer distance $i$
2:Each node in $R_{i}$ is assigned exactly $i$ interior neighbors
3:Initialize a dictionary $assigned\_neighbors$ mapping each node to an empty set
4:Let $n\leftarrow|R_{i}|$
5:for each node $u\in R_{i}$ do
6:     $assigned\leftarrow 0$
7:     for $step=1$ to $n-1$ do
8:         $v\leftarrow R_{i}[(\text{index}(u)+step)\bmod n]$
9:         if $v\neq u\And v\notin assigned\_neighbors[u]$ then
10:              Add $v$ to $assigned\_neighbors[u]$
11:              $assigned\leftarrow assigned+1$
12:         end if
13:         if $assigned=i$ then
14:              break
15:         end if
16:     end for
17:end forreturn $assigned\_neighbors$
Algorithm 2 Given a rooted neighborhood $R_{i}$, this algorithm assigns each node to participate in exactly $i$ interior cycles, simulating repeated interior degree mappings for each level. This construction ensures every node maps to interior neighbors via multiple cycle layers, supporting the interior degree doubling condition.
@@
Even with limitations on 2-cycles, such a mapping algorithm will allow for the utilization of combinatorial space in $R_{i}$. We can actually look at this as we begin in the rooted neighborhood $R_{i}$.
@@
This results in $i!$ possible candidate paths (modulo orientation constraints). This number of such derangements still grows factorially at roughly $\frac{n!}{e}$. This indicates that we still have an abundance of cycle permutations with the 2-cycle restrictions.
The structure of these rooted neighborhoods forms a permutation on the set $R_{i}$, which decomposes into disjoint cycles. Since oriented graphs prohibit 2-cycles, each of these cycles must have a length $k\geq 3$. A key result from group theory (Lagrange’s Theorem) tells us that the order of any subgroup divides the order of the full symmetric group on $R_{i}$, $S_{|R_{i}|}$. These permutations decompose into disjoint cycles whose lengths divide $|R_{i}|$, and the cycles are structure-preserving under composition. Because Lagrange’s theorem partitions $R_{i}$, into subsets that divide $|R_{i}|$, we know that any remaining permutations will always divide the order of the group and no nodes will ever be left without a cycle.
Thus, each node is mapped to a unique successor in its cycle. Therefore, in every cycle of length $k\geq 3$, each node has out-degree 1 and maps to another node in the same cycle, producing a single out-edge within $R_{i}$. The node it maps to will also have a unique out-neighbor since cycles do not repeat nodes. This yields a second-level interior neighbor.
Consider any node $v$ within one of these cycles of length $k\geq 3$. By definition of a cycle within $R_{i}$, $v$ has exactly one outgoing edge $v\to u$ where $u\in R_{i}$. Therefore, $|N^{+}_{int}(v)|=1$. Furthermore, since $k\geq 3$, $u$ must also have exactly one outgoing edge $u\to w$ where $w\in R_{i}$ and $w\neq v$. Thus, $w$ is a unique second-level interior neighbor, implying $|N^{++}_{int}(v)|=1$. Since for every node $v$ in any such rooted neighborhood $R_{i}$, we have
$|N^{+}_{\text{int}}(v)|=1$
and
$|N^{++}_{\text{int}}(v)|=1,$
the condition
$|N^{++}_{\text{int}}(v)|\geq|N^{+}_{\text{int}}(v)|$
is satisfied.
@@
This result is also aligned with Lemma 4.1 (Minimum Out-Degree $<3$) and related results from [20], confirming that low out-degree nodes within cycles tend to satisfy the degree doubling property. ∎
@@
The positive answer to interior degree-doubling is helpful to the construction of our algorithm. In the absence of back arcs, we can split a node’s degree into interior and exterior degrees. Knowing then that when a node is forced to take upon $i$ cycles in the rooted neighborhood $R_{i}$, that node’s interior degree will have to double means that if that node’s exterior degree also doubles, then that node’s overall degree will double. This gives light to our overall strategy. We will first prove a generalization of the interior load balancing lemma. To do this, though, we need to first prove that the exterior neighbors decrease as we traverse a path. We will assume that all nodes are bound by the Decreasing Neighborhood Sequence Property (DNSP). Then if any node has both its interior and exterior degree double, its overall degree will double, which will lead to a contradiction of the DNSP.
@@
Lemma 6.4 (Interior Degrees Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the rooted neighborhoods as an oriented graph and searched for a node whose degree doubled inside that rooted neighborhood. The proof uses abstract algebra: we can always construct the necessary cycles in the rooted neighborhood $R_{i}$ to make a node’s degree double. We still need to show that these cycles generalize to arbitrary rooted neighborhoods, which is an important part of this result.
@@
This also shows the reach that SSNC has across mathematics. Among the published literature there is not much work in the field of abstract algebra connected to the SSNC. However, given the importance of cycles in keeping nodes from having their degrees double, we are able to see that we are always able to find a representation that maps nodes to interior degrees in that many disjoint cycles. The disjoint cycle is what allows the interior degrees to double, because each cycle has nodes of degree one on the cycle itself.
@@
By finding a positive answer to this localized version of the conjecture, and not just a positive answer but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture. Not only that, but it helps us to dissect the problem down into two components: interior degree and exterior degree. If we know that the interior degree of every node inside the rooted neighborhoods doubles, then we can turn our attention to the exterior degrees.
@@
We will then be able to introduce a traversal algorithm that explores these exterior degrees. It will systematically search the nodes of $G$ for a node whose exterior degree doubles. That is, we will question whose second (exterior) neighborhood is at least as large as its first. Lemma 6.6 (Decreasing Exteriors Lemma) will show that these exterior neighborhoods must shrink (decrease) at each step. We will then use Lemma 6.6 to show that we can generalize the interior load balancing lemma. Once we have proved these two major results, it will go a long way towards proving our contradiction.